OpenStudy (anonymous):

using integration by substitution what is the integral of 2x+3/x^2+3x-5

myininaya (myininaya):

Try letting u=the bottom

OpenStudy (anonymous):

Let u = x^2 + 3x - 5 then du = (2x+3)dx and you can solve it. Remember that integral (1/u)du = ln|u|.

OpenStudy (anonymous):